Let H be a Hilbert space.

(i) Let S ⊆ H be any non-empty subset of S. Show that S ^⊥ is a subspace of H.


(ii) Let L ⊆ H be a linear manifold. Show that L^ ⊥⊥ = L


1
Expert's answer
2021-02-17T02:54:46-0500

1. Let S ⊆ H be any non-empty subset of S. "S^\perp=\{x\in H: \forall s\in S \langle x,s\rangle=0\}"

Let "x,y\in S^\perp". Then for any "s\in S" we have "\langle x+y,s\rangle=\langle x,s\rangle+\langle y,s\rangle=0+0=0" and for any scalar "\lambda" we have "\langle\lambda x,s\rangle=\lambda\langle x,s\rangle=0". Therefore, "S^\perp" is a linear.

If "\{x_n\}_{n=1}^{+\infty}\subset S^\perp" converges to "x\in H", then for any "s\in S" we have "\langle x,s\rangle=\lim_{n\to+\infty}\langle x_n,s\rangle=0".

This implies that "S^\perp" is a closed subspace of H.


2. Let L ⊆ H be a linear manifold (i.e. closed linear subspace) of H. For all "x\in L" and "y\in L^\perp" we have "\langle x,y\rangle=0", therefore, by the definition, "L\subset L^{\perp\perp}".

If "x\notin L" then let "x_L" be an orthogonal projection of to L. It can be obtained as follows.

L is a closed linear subspace of H, therefore L is a Hilbert space. Let "\{v_1,v_2,...\}" - the orthonormal basis of L, "c_n=\langle x, v_n\rangle", "x_n=\sum_{k=0}^{n}c_kv_k". Then if k<n then "\langle x_n, v_k\rangle=\langle x, v_k\rangle" and "\langle x-x_n, v_k\rangle=0". From the Pythagorean theorem "|x|^2=|(x-x_n)+x_n|^2=|x-x_n|^2+|x_n|^2", from where "|x_n|^2\leq |x|^2" . From the other side, "|x_n|^2=\sum_{k=0}^{n}|c_k|^2" , therefore the partial sums of the series "\sum_{k=0}^{+\infty}|c_k|^2" are bounded and the sequence "x_n" is fundamental in L, hence, it converges to some vector "x_L\in L". Since "\langle x-x_n, v_k\rangle=0" , then "\langle x-x_L, v_k\rangle=\lim_{n\to+\infty}\langle x-x_n, v_k\rangle=0" for any k. Then "x-x_L\in L^\perp" and "\langle x-x_L,x\rangle=\langle x-x_L,x\rangle-\langle x-x_L,x_L\rangle=\langle x-x_L,x-x_L\rangle=|x-x_L|^2>0"

This shows that "x\notin L^{\perp\perp}" and therefore, "L=L^{\perp\perp}".